# Leave-One-Out Neighborhood Smoothing for Graphons: Berry-Esseen Bounds, Confidence Intervals, and Honest Tuning

### v3

@@ Abstract
Estimating latent edge probabilities in network models often relies on aggregating information from structurally similar nodes. However, using the same adjacency matrix to both determine node similarities and estimate probabilities induces complex statistical dependencies. This coupling invalidates standard inferential tools, such as concentration inequalities and central limit theorems, leaving a critical gap in entrywise uncertainty quantification for networks. To resolve this, we introduce a leave-one-out (LOO) decoupling strategy. When predicting a target edge, the local neighborhood is constructed from a reduced adjacency matrix with the target row and column entirely removed. This structural modification-rather than the specific smoothing mechanism-is the core innovation: it restores exact conditional independence between the data-driven neighborhood selection and the observed edges being aggregated. Leveraging this restored independence, we derive variance-adaptive empirical Bernstein bounds and establish uniform Berry-Esseen normal approximations with explicit rates. These results enable the construction of valid finite-sample and asymptotic confidence intervals for individual edge probabilities. Furthermore, the LOO architecture naturally admits an honest cross-validation scheme for tuning parameter selection, providing a complete and rigorous inferential framework while preserving strong mean-squared error performance.
@@ 1 Introduction
Graphon models have emerged as the canonical nonparametric framework for analyzing complex network data, providing a flexible representation of latent edge probabilities through a measurable kernel on the unit square (Borgs et al., 2007; Lovász, 2012). Over the past decade, a rich literature has developed around estimating the underlying probability matrix $P$ or the graphon $h$ from a single observed adjacency matrix $A$. Early foundational work focused heavily on the Stochastic Block Model (SBM) and its variants, utilizing modularity maximization, maximum likelihood, and spectral clustering to recover block probabilities (Bickel and Chen, 2009; Abbe, 2017; Rohe et al., 2011). As the field moved toward continuous, nonparametric graphons, modern estimators-including universal singular value decomposition (Chatterjee, 2015), sorting-based smoothing (Chan and Airoldi, 2014), and block-approximations (Gao et al., 2015)-have been proven to reliably achieve minimax-optimal rates for point estimation.
@@
This decoupling is the central theoretical lever of the paper. By strictly separating the data used for neighborhood selection from the data used for estimation, we restore exact conditional independence among the centered summands. This topological decoupling bridges the gap between network estimation and classical empirical process theory. Specifically, under standard piecewise Lipschitz assumptions on the graphon, we leverage this restored independence to establish three main contributions:
@@
    Asymptotic Normality: We establish uniform Berry-Esseen bounds with explicit convergence rates for the normalized estimation error, enabling the construction of valid normal-approximation confidence intervals for individual edge probabilities.
@@
    Honest Cross-Validation: The LOO structure naturally supports an unbiased cross-validation scheme for tuning the neighborhood size $h_{n}$.inequality.
Crucially, these inferential guarantees do not come at the cost of statistical efficiency. We prove that the proposed LOO estimator preserves the optimal row-wise mean-squared error rates of classical neighborhood smoothing.
@@ 2 Model and Leave-One-Out Neighborhood Smoothing
The edge probabilities satisfy $P_{ij}=h(\xi_{i},\xi_{j})$ with $\xi_{1},\dots,\xi_{n}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Unif}[0,1]$. The graphon $h:[0,1]^{2}\to[0,1]$ is piecewise Lipschitz in each argument with Lipschitz constant $L$ on a fixed finite rectangular partition of $0,1]^{2}$.
@@
$\eta\leq h(u,v)\leq 1-\eta\qquad\text{for all }(u,v)\in[0,1]^{2}.$
@@
Assumption 2.2 requires that there exists $\eta\in(0,1/2)$ such that $\eta\leq h(u,v)\leq 1-\eta$ for all $(u,v)\in[0,1]^{2}$. This nondegeneracy condition guarantees a uniformly nonvanishing conditional variance and is used to apply Berry-Esseen and variance‑adaptive concentration bounds. It excludes extremely sparse regimes where $P_{ij}\to 0$ uniformly; extending our results to such regimes requires variance re‑normalization and additional control of rare events, which we leave to future work. Empirically, moderate thinning of edges does not invalidate the procedure (see Section 5).
@@
For concreteness we work in the canonical (dense) graphon regime $P_{ij}=h(\xi_{i},\xi_{j})$ with $h$ fixed as in Assumption 2.1. In this regime the expected degree satisfies $\mathbb{E}\deg(i)=\sum_{j=1}^{n}P_{ij}\asymp n$, so local neighborhoods contain many informative two‑hop paths and several concentration steps simplify. All main results below are stated under Assumptions 2.1 and 2.2 together with this dense regime convention. Extensions to explicit sparsity scalings $P_{ij}=\rho_{n}h(\xi_{i},\xi_{j})$ are discussed briefly in Section 6.
@@ 2.1 LOO neighborhoods and estimator
Our estimator is based on an adaptive neighborhood constructed through a leave-one-out (LOO) modification of the Zhang-Levina-Zhu (ZLZ) distance. For $i\neq j$, define the empirical second‑order (ZLZ) distance (Zhang et al., 2017)
$d_{ij}:=\max_{k\neq i,j}\Big|\big(A^{2}/n\big)_{ik}-\big(A^{2}/n\big)_{jk}\Big|.$ (2.2)
This is the sup‑norm discrepancy between the $i$th and $j$th rows of the normalized two‑hop matrix $A^{2}/n$, measuring the similarity of their two‑hop neighborhoods.
For estimation of a fixed entry $(i,j)$, remove node $j$ and form the reduced adjacency matrix $A^{(-j)}$ obtained by deleting the $j$th row and column. Define the two-hop matrix
$M^{(-j)}=\frac{1}{n-1}A^{(-j)}A^{(-j)}.$ (2.3)
For node $i$ and node $k\neq i,j$, define the ZLZ distance as follows:
@@
$d_{i,(1)}^{(-j)}\leq\cdots\leq d_{i,(n-2)}^{(-j)},$
denote the order statistics of all nodes except $j$ to node $i$, namely $\{d_{ik}^{(-j)}:k\neq i,j\}$. Define the leave-one-out neighborhood of node $i$ by
@@
that is, the $h_{n}$ nearest neighbors of $i$ according to the ZLZ two-hop distance computed on the reduced graph.
@@
The classical neighborhood smoother of Zhang et al. (2017) uses $\mathcal{N}_{i}$ computed from the full adjacency matrix $A$ to estimate the entire $i$th row of $P$. Because $\mathcal{N}_{i}$ depends on all columns of $A$, including column $j$, the weights used to average $\{A_{kj}:k\in\mathcal{N}_{i}\}$ depend on the very summands being averaged. Hence the centered terms $A_{kj}-P_{kj}$ are not independent conditional on the neighborhoods, preventing the direct application of variance‑adaptive concentration or Berry-Esseen theory.
$\tilde{P}_{ij}:=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj},\qquad\widehat{P}_{ij}:=\frac{\tilde{P}_{ij}+\tilde{P}_{ji}}{2}.$ (2.6)
@@
Condition on the latent positions $(\xi_{1},\dots,\xi_{n})$ so that $P$ is deterministic with $P_{kl}=h(\xi_{k},\xi_{\ell})$. Under the model (2.1), the edges $\{A_{k\ell}:1\leq k<\ell\leq n\}$ are independent Bernoulli random variables with parameters $P_{k\ell}$.
@@ 3 Consistency
In this section we quantify the deviation of the LOO estimator $\tilde{P}_{ij}$ from the target $P_{ij}$. The analysis separates the estimation error into a variance component, arising from the randomness of observed edges given $P$, and a bias component, arising from approximating $P_{ij}$ by a local average over a neighborhood of $i$. We begin with a decomposition that will be used throughout the section. For $i\neq j$ write For $i\neq j$, write
$\Delta_{ij}:=\tilde{P}_{ij}-P_{ij}=U_{ij}+B_{ij},$
@@
$U_{ij}^{(-j)}$ $:=\tilde{P}_{ij}-\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj}),$ (3.1)
$B_{ij}^{(-j)}$ $:=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}-P_{ij}.$ (3.2)
By Proposition 2.7, conditional on $\mathcal{N}_{i}^{(-j)}$ and $P$ the summands in $U_{ij}^{(-j)}$ are independent, mean zero and bounded, so $U_{ij}^{(-j)}$ behaves like an average of $h$ independent observations. The term $B_{ij}^{(-j)}$ is a deterministic (given $P$ and $\mathcal{N}_{i}^{(-j)}$) smoothness bias that can be controlled using Assumption 2.1 and the neighborhood properties in Section 2.
@@ 3.1 Neighborhood bias
Let $A\in\{0,1\}^{n\times n}$ be the adjacency matrix of an undirected simple graph with $A_{ii}=0$. For each $j\in[n]$, let $A^{(-j)}$ be the $(n-1)\times(n-1)$ adjacency matrix obtained by deleting the $j$th row and column (i.e., removing vertex $j$). Define the normalized two-hop matrices
@@
For each $j$, embed $M^{(-j)}$ into an $n\times n$ matrix by placing it on the index set $n]\setminus\{j\}$ and setting entries involving index $j$ to zero. Then, deterministically,
$\sup_{1\leq j\leq n}\ \|M-M^{(-j)}\|_{\max}\ \leq\ \frac{3}{n},$
where $\|B\|_{\max}:=\max_{1\leq i,\ell\leq n}|B_{i\ell}|$ denotes the entrywise max norm. In particular, for every $j\in[n]$ and all $i,\ell\in[n]\setminus\{j\}$,
$\big|M_{i\ell}-M^{(-j)}_{i\ell}\big|\leq\frac{3}{n}.$
@@
Thus the bound holds for every $j$ (in fact with constant $2$), and taking the supremum over $j$ yields $\|M-M^{(-j)}\|_{\max,[n]\setminus\{j\}}:=\max_{i,\ell\in[n]\setminus\{j\}}|M_{i\ell}-M^{(-j)}_{i\ell}|\leq 2/n$. ∎
Lemma 3.2 (Two-hop concentration is preserved under leave-one-out deletion).
Let $A\in\{0,1\}^{n\times n}$ be generated from the graphon model (2.1) with probability matrix $P$. Define
$M:=\frac{1}{n}A^{2},\qquad M^{\star}:=\mathbb{E}\!\left[M\,\middle|\,P\right],$
and for each $j\in[n]$ let $A^{(-j)}$ be the adjacency matrix obtained by deleting the $j$th row and column. Define the leave-one-out two-hop matrices
$M^{(-j)}:=\frac{1}{n-1}\big(A^{(-j)}\big)^{2},\qquad M^{(-j)\star}:=\mathbb{E}\!\left[M^{(-j)}\,\middle|\,P\right].$
Embed $M^{(-j)}$ and $M^{(-j)\star}$ into $n\times n$ matrices on the index set $n]\setminus\{j\}$ (and set entries involving $j$ to zero).
Then, for every $n$,
@@
for a universal constant $C>0$. In particular, if
$\|M-M^{\star}\|_{\max}=\mathcal{O}_{\mathbb{P}}\!\Big(\sqrt{\tfrac{\log n}{n}}\Big),$
(as established for the ZLZ two-hop matrix in Zhang et al. (2017) and its supplementary material), then
$\sup_{1\leq j\leq n}\ \big\|M^{(-j)}-M^{(-j)\star}\big\|_{\max}=\mathcal{O}_{\mathbb{P}}\!\Big(\sqrt{\tfrac{\log n}{n}}\Big).$
@@
Fix $j\in[n]$. By the triangle inequality,
$\big\|M^{(-j)}-M^{(-j)\star}\big\|_{\max}$ $\leq\big\|M^{(-j)}-M\big\|_{\max}+\big\|M-M^{\star}\big\|_{\max}+\big\|M^{\star}-M^{(-j)\star}\big\|_{\max}.$
The first term is deterministic and satisfies
$\sup_{1\leq j\leq n}\|M^{(-j)}-M\|_{\max}\leq\frac{C_{1}}{n}$
by Lemma 3.1. For the last term, note that
$M^{\star}=\mathbb{E}[M\mid P],\qquad M^{(-j)\star}=\mathbb{E}[M^{(-j)}\mid P],$
so Jensen’s inequality and Lemma 3.1 give
$\sup_{1\leq j\leq n}\|M^{\star}-M^{(-j)\star}\|_{\max}=\sup_{1\leq j\leq n}\big\|\mathbb{E}[M-M^{(-j)}\mid P]\big\|_{\max}\leq\sup_{1\leq j\leq n}\mathbb{E}\big[\|M-M^{(-j)}\|_{\max}\mid P\big]\leq\frac{C_{1}}{n}.$
Combining the bounds yields (3.3) with $C=2C_{1}$. Finally, since $n^{-1}=o\!\big(\sqrt{(\log n)/n}\big)$, the stated rate transfer follows immediately. ∎
Proposition 3.3 (ZLZ neighborhood smoothing: row-risk (black box)).
Work under the assumptions of Zhang, Levina, and Zhu (2017, Biometrika) and choose the bandwidth $h\asymp\sqrt{n\log n}$ as in their Theorem 1. Let $N_{i}$ be the ZLZ neighborhood of node $i$ constructed from the two-hop matrix $M=A^{2}/n$ via their distance and thresholding rule, and define the ZLZ estimator
$\tilde{P}^{\mathrm{ZLZ}}_{ij}:=\frac{1}{|N_{i}|}\sum_{k\in N_{i}}A_{kj}.$
Then there exist constants $C,c>0$ such that with probability at least $1-n^{-c}$, uniformly over $i$,
$\frac{1}{n}\sum_{j\neq i}\big(\tilde{P}^{\mathrm{ZLZ}}_{ij}-P_{ij}\big)^{2}\;\leq\;C\Big(\frac{\log n}{n}\Big)^{1/2}.$
@@
This is the row-wise conclusion proved for $\tilde{P}_{ij}$ in Zhang et al. (2017), using their supplementary decomposition into $J_{1}(i,j)$ and $J_{2}(i,j)$ (equation 1 in the supplementary material of Zhang et al. (2017)) and the two-hop concentration input. ∎
Corollary 3.4 (Leave-one-out ZLZ row-wise risk + $(U_{ij},B_{ij})$ decomposition).
Fix $j\in[n]$. Let $A^{(-j)}$ be the adjacency matrix with row/column $j$ deleted, and let $\widetilde{P}^{(-j)}$ denote the ZLZ neighborhood-smoothing estimator computed from $A^{(-j)}$ (with neighborhoods $N_{i}^{(-j)}$ defined exactly as in ZLZ but using the two-hop matrix $(A^{(-j)})^{2}/(n-1)$). For $i\neq j$ and $k\in N_{i}^{(-j)}$, define
$U^{(-j)}_{ij}:=\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}(A_{kj}-P_{kj}),\qquad B^{(-j)}_{ij}:=\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}(P_{kj}-P_{ij}),$

### v4

@@ Abstract
Estimating latent edge probabilities in network models often relies on aggregating information from structurally similar nodes. However, using the same adjacency matrix to both determine node similarities and estimate probabilities induces complex statistical dependencies. This coupling invalidates standard inferential tools, such as concentration inequalities and central limit theorems, leaving a critical gap in entrywise uncertainty quantification for networks. To resolve this, we introduce a leave-one-out (LOO) decoupling strategy. When predicting a target edge, the local neighborhood is constructed from a reduced adjacency matrix with the target row and column entirely removed. This structural modification-rather than the specific smoothing mechanism-is the core innovation: it restores exact conditional independence between the data-driven neighborhood selection and the observed edges being aggregated. Leveraging this restored independence, we derive variance-adaptive empirical Bernstein bounds and establish uniform Berry-Esseen normal approximations with explicit rates. These results enable finite-sample, variance-adaptive confidence intervals for the localized neighborhood mean and entrywise uncertainty quantification for $P_{ij}$ on a high-probability interior (good) set up to an explicit bias term. We also establish a uniform Berry–Esseen approximation for the variance component (normal‑approximation intervals use a heuristic variance plug‑in in this version).
Furthermore, the LOO architecture naturally admits an honest cross-validation scheme for tuning parameter selection, providing a complete and rigorous inferential framework while preserving strong mean-squared error performance.
@@ 1 Introduction
Graphon models have emerged as the canonical nonparametric framework for analyzing complex network data, providing a flexible representation of latent edge probabilities through a measurable kernel on the unit square (Borgs et al., 2008; Lovász, 2012). Over the past decade, a rich literature has developed around estimating the underlying probability matrix $P$ or the graphon $f$ from a single observed adjacency matrix $A$. Early foundational work focused heavily on the Stochastic Block Model (SBM) and its variants, utilizing modularity maximization, maximum likelihood, and spectral clustering to recover block probabilities (Bickel and Chen, 2009; Abbe, 2017; Rohe et al., 2011). As the field moved toward continuous, nonparametric graphons, modern estimators-including universal singular value decomposition (Chatterjee, 2015), sorting-based smoothing (Chan and Airoldi, 2014), and block-approximations (Gao et al., 2015)-have been proven to reliably achieve minimax-optimal rates for point estimation.
@@
This decoupling is the central theoretical lever of the paper. By strictly separating the data used for neighborhood selection from the data used for estimation, we restore exact conditional independence among the centered summands. This topological decoupling permits the application of classical empirical process tools to graphon neighborhood smoothing. Specifically, under standard piecewise Lipschitz assumptions on the graphon, we leverage this restored independence to establish three main contributions:
@@
    Asymptotic Normality: We establish a uniform Berry–Esseen bound with explicit rates for the normalized variance component; normal‑approximation intervals in this version use a heuristic variance estimate and incur an additional variance‑estimation error term.
@@
    Honest Cross-Validation: The LOO structure naturally supports an unbiased cross-validation scheme for tuning the neighborhood size $h_{n}$.inequality. Here, honest cross-validation means that it does not reuse held-out edges in predictor construction, ensuring valid inference even after data-driven tuning.
Crucially, these inferential guarantees preserve the optimal rowwise mean-squared error rate of classical neighborhood smoothing.
@@ 2 Model and Leave-One-Out Neighborhood Smoothing
The edge probabilities satisfy $P_{ij}=f(\xi_{i},\xi_{j})$ with $\xi_{1},\dots,\xi_{n}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Unif}[0,1]$. The graphon $f:[0,1]^{2}\to[0,1]$ is piecewise Lipschitz in each argument with Lipschitz constant $L$ on a fixed finite rectangular partition of $0,1]^{2}$.
@@
$\eta\leq f(u,v)\leq 1-\eta\qquad\text{for all }(u,v)\in[0,1]^{2}.$
Assumption 2.2 places the model in the dense graphon regime, where edge probabilities are uniformly bounded away from zero and one. This excludes sparse scalings of the form $P_{ij}=\rho_{n}W(\xi_{i},\xi_{j})$ with $\rho_{n}\to 0$. Our Berry-Esseen and empirical Bernstein arguments rely critically on variance bounded away from zero. Extending the present theory to sparse or degree-corrected regimes would require variance renormalization and additional control of small-probability events, and is left for future work.
@@
Assumption 2.2 requires that there exists $\eta\in(0,1/2)$ such that $\eta\leq f(u,v)\leq 1-\eta$ for all $(u,v)\in[0,1]^{2}$. This nondegeneracy condition guarantees a uniformly nonvanishing conditional variance and is used to apply Berry-Esseen and variance‑adaptive concentration bounds. It excludes extremely sparse regimes where $P_{ij}\to 0$ uniformly; extending our results to such regimes requires variance re‑normalization and additional control of rare events, which we leave to future work. Empirically, moderate thinning of edges does not invalidate the procedure (see Section 5).
@@
For concreteness we work in the canonical (dense) graphon regime $P_{ij}=f(\xi_{i},\xi_{j})$ with $f$ fixed as in Assumption 2.1. In this regime the expected degree satisfies $\mathbb{E}\deg(i)=\sum_{j=1}^{n}P_{ij}\asymp n$, so local neighborhoods contain many informative two‑hop paths and several concentration steps simplify. All main results below are stated under Assumptions 2.1 and 2.2 together with this dense regime convention. Extensions to explicit sparsity scalings $P_{ij}=\rho_{n}f(\xi_{i},\xi_{j})$ are discussed briefly in Section 6.
@@ 2.1 LOO neighborhoods and estimator
Consider the set of nodes $V=\{1,\dots,n\}$. Fix $j\in[n]$ and let $A^{(-j)}$ be the $(n-1)\times(n-1)$ adjacency matrix obtained by deleting row and column $j$ from $A$. Define the two-hop matrices
$M:=\frac{1}{n}A^{2},\qquad M^{(-j)}:=\frac{1}{n-1}\big(A^{(-j)}\big)^{2}.$
For distinct $i,k\in V\setminus\{j\}$, define the leave-one-out (LOO)-distance
@@
$d^{(-j)}_{i,(1)}\leq\cdots\leq d^{(-j)}_{i,(n-2)}$
denote the order statistics of the set
$\bigl\{d^{(-j)}(i,k):k\in V\setminus\{i,j\}\bigr\}.$
Define the leave-one-out neighborhood of node $i$ by
@@
where $h_{n-1}$ is the ZLZ bandwidth parameter with $n$ replaced by n1n-1. that is, the $h_{n}$ nearest neighbors of $i$ according to the ZLZ two-hop distance computed on the reduced graph.
$\tilde{P}_{ij}:=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj},\qquad\widehat{P}_{ij}:=\frac{\tilde{P}_{ij}+\tilde{P}_{ji}}{2}.$ (2.4)
@@
The classical neighborhood smoother of Zhang et al. (2017) uses $\mathcal{N}_{i}$ computed from the full adjacency matrix $A$ to estimate the entire $i$th row of $P$. Define the distance function $d:V\times V\to\mathbb{R}_{+}$ by
$d_{ZLZ}(i,j):=\begin{cases}0,&\text{for }i=j,\\ \max_{k\neq i,j}\Big|\big(A^{2}/n\big)_{ik}-\big(A^{2}/n\big)_{jk}\Big|,&\text{for }i\neq j.\end{cases}$ (2.5)
This is the sup‑norm discrepancy between the $i$th and $j$th rows of the normalized two‑hop matrix $A^{2}/n$, measuring the similarity of their two‑hop neighborhoods. Therefore the ZLZ neighborhood $\mathcal{N}_{i}$ consists of the $h_{n}$ nodes with the smallest ZLZ distance to $i$. Zhang et al. (2017) defined
$\mathcal{N}_{i}=\{k\in V\setminus\{i\}:d_{ZLZ}(i,k)\leq q_{h}(\{D\}),\quad D=\{d_{ZLZ}(i,k):k\in V\setminus\{i\}\}\},$ (2.6)
and $q_{h}(\{D\})$ is the $h_{n}$-th order statistic of the set $\{d_{ZLZ}(i,k):k\in V\setminus\{i\}\}$. The ZLZ estimator is then defined by
$\widetilde{P}^{\mathrm{ZLZ}}_{ij}:=\frac{1}{|\mathcal{N}_{i}|}\sum_{k\in\mathcal{N}_{i}}A_{kj}.$
Because $\mathcal{N}_{i}$ depends on all columns of $A$, including column $j$, the weights used to average $\{A_{kj}:k\in\mathcal{N}_{i}\}$ depend on the very summands being averaged. Hence the centered terms $A_{kj}-P_{kj}$ are not independent conditional on the neighborhoods, preventing the direct application of variance‑adaptive concentration or Berry-Esseen theory.
@@
The leave-one-out construction does not alter the smoothing mechanism itself; it alters only the dependency structure. By removing the target column prior to neighborhood selection, the summands in the neighborhood average become exactly independent conditional on $P$. This structural decoupling is the key step that allows classical empirical process tools to be applied without resampling or sample splitting.
@@
Condition on the latent positions $(\xi_{1},\dots,\xi_{n})$ so that $P$ is deterministic with $P_{kl}=f(\xi_{k},\xi_{\ell})$. Under the model (2.1), the edges $\{A_{k\ell}:1\leq k<\ell\leq n\}$ are independent Bernoulli random variables with parameters $P_{k\ell}$.
@@ 3 Consistency
Our objective is entrywise inference for the finite-sample edge probability matrix $P=(P_{ij})_{1\leq i,j\leq n}$, where $P_{ij}=W(\xi_{i},\xi_{j})$. All probability statements in Sections 3 and 4 are conditional on the realized latent positions $(\xi_{1},\dots,\xi_{n})$, equivalently on $P$. Thus our results provide conditional finite-sample and asymptotic inference for $P_{ij}$ rather than for the graphon $W$ itself. In this section we quantify the deviation of the LOO estimator $\tilde{P}_{ij}$ from the target $P_{ij}$. The analysis separates the estimation error into a variance component, arising from the randomness of observed edges given $P$, and a bias component, arising from approximating $P_{ij}$ by a local average over a neighborhood of $i$. We begin with a decomposition that will be used throughout the section. For $i\neq j$ write For $i\neq j$, write
$\Delta_{ij}:=\tilde{P}_{ij}-P_{ij}=U_{ij}^{(-j)}+B_{ij}^{(-j)},$ (3.1)
@@
$U_{ij}^{(-j)}$ $:=\tilde{P}_{ij}-\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj}),$ (3.2)
$B_{ij}^{(-j)}$ $:=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}-P_{ij}.$ (3.3)
By Proposition 2.7, conditional on $\mathcal{N}_{i}^{(-j)}$ and $P$ the summands in $U_{ij}^{(-j)}$ are independent, mean zero and bounded, so $U_{ij}^{(-j)}$ behaves like an average of $h_{n}$ independent observations. The term $B_{ij}^{(-j)}$ is a deterministic (given $P$ and $\mathcal{N}_{i}^{(-j)}$) smoothness bias that can be controlled using Assumption 2.1 and the neighborhood properties in Section 2. The following two subsections analyze these two components separately.
@@ 3.1 Neighborhood bias
Let $A\in\{0,1\}^{n\times n}$ be the adjacency matrix of an undirected simple graph with $A_{ii}=0$. For each $j\in[n]$, let $A^{(-j)}$ be the $(n-1)\times(n-1)$ adjacency matrix obtained by deleting the $j$th row and column. Define the normalized two-hop matrices
@@
To compare $M$ and $M^{(-j)}$, embed $M^{(-j)}$ into an $n\times n$ matrix $M^{(-j)}_{\mathrm{embed}}$ by
$M^{(-j)}_{\mathrm{embed}}(i,\ell)=\begin{cases}M^{(-j)}_{i\ell},&i,\ell\neq j,\\ M_{i\ell},&\text{if }i=j\text{ or }\ell=j.\end{cases}$
Thus $M^{(-j)}_{\mathrm{embed}}$ and $M$ have the same dimensions and agree on the $j$th row and column.
Then, deterministically,
$\sup_{1\leq j\leq n}\ \|M-M^{(-j)}_{\mathrm{embed}}\|_{\max}\;\leq\;\frac{2}{n},$
where $\|B\|_{\max}:=\max_{1\leq i,\ell\leq n}|B_{i\ell}|$. In particular, for all $j\in[n]$ and all $i,\ell\in[n]\setminus\{j\}$,
$|M_{i\ell}-M^{(-j)}_{i\ell}|\leq\frac{2}{n}.$
@@
For $i=j$ or $\ell=j$, we have $M^{(-j)}_{\mathrm{embed}}(i,\ell)=M_{i\ell}$ by definition, so the difference is zero. Hence
$\|M-M^{(-j)}_{\mathrm{embed}}\|_{\max}\leq\frac{2}{n}$
for each $j$, and taking the supremum over $j$ preserves the bound. ∎
Lemma 3.2 (Stability of expected two-hop matrices under deletion).
Let $A$ and $A^{(-j)}$ be as in Lemma 3.1, and define
$M=\frac{1}{n}A^{2},\qquad M^{\star}=\mathbb{E}[M\mid P],\qquad M^{(-j)}=\frac{1}{n-1}(A^{(-j)})^{2},\qquad M^{(-j)\star}=\mathbb{E}[M^{(-j)}\mid P].$
Embed $M^{(-j)}$ and $M^{(-j)\star}$ into $n\times n$ matrices as in Lemma 3.1. Then, for all $n$,
@@
for a universal constant $C>0$.
@@
Fix $j\in[n]$. Let $M^{(-j)}_{\mathrm{embed}}$ and $M^{(-j)\star}_{\mathrm{embed}}$ be the $n\times n$ embeddings of $M^{(-j)}$ and $M^{(-j)\star}$ defined as in Lemma 3.1. Then
$\|M^{(-j)}_{\mathrm{embed}}-M^{(-j)\star}_{\mathrm{embed}}\|_{\max}\leq\|M^{(-j)}_{\mathrm{embed}}-M\|_{\max}+\|M-M^{\star}\|_{\max}+\|M^{\star}-M^{(-j)\star}_{\mathrm{embed}}\|_{\max}.$
By Lemma 3.1,
$\sup_{1\leq j\leq n}\|M^{(-j)}_{\mathrm{embed}}-M\|_{\max}\leq\frac{2}{n}.$
Next, since
$M^{\star}=\mathbb{E}[M\mid P],\qquad M^{(-j)\star}_{\mathrm{embed}}=\mathbb{E}[M^{(-j)}_{\mathrm{embed}}\mid P],$
we have
$M^{\star}-M^{(-j)\star}_{\mathrm{embed}}=\mathbb{E}[M-M^{(-j)}_{\mathrm{embed}}\mid P],$
and thus, by Jensen’s inequality and Lemma 3.1,
$\|M^{\star}-M^{(-j)\star}_{\mathrm{embed}}\|_{\max}\leq\mathbb{E}\big[\|M-M^{(-j)}_{\mathrm{embed}}\|_{\max}\mid P\big]\leq\frac{2}{n}.$
Combining the bounds,
$\|M^{(-j)}_{\mathrm{embed}}-M^{(-j)\star}_{\mathrm{embed}}\|_{\max}\leq\|M-M^{\star}\|_{\max}+\frac{4}{n}.$
Taking the supremum over $j$ yields (3.4). ∎
Proposition 3.3 (Theorem 1 of Zhang et al. (2017): Row-wise risk).
Work under the assumptions of Zhang et al. (2017) and choose the ZLZ quantile cutoff parameter $h_{\mathrm{ZLZ}}=C_{0}(\log n/n)^{1/2}$, so that the resulting neighborhood sizes satisfy $|N_{i}|\geq(n\log n)^{1/2}$ (Lemma 2(1) in their supplementary material). Let $N_{i}$ be the ZLZ neighborhood of node $i$ constructed from the two-hop matrix $M=A^{2}/n$ via their distance-and-quantile rule, and define
$\widetilde{P}^{\mathrm{ZLZ}}_{ij}:=\frac{1}{|N_{i}|}\sum_{k\in N_{i}}A_{kj}.$
Then there exist constants $C,c>0$ such that with probability at least $1-Cn^{-c}$, uniformly over all $i\in[n]$,
$\frac{1}{n}\sum_{j=1}^{n}\big(\widetilde{P}^{\mathrm{ZLZ}}_{ij}-P_{ij}\big)^{2}\;\leq\;C\Big(\frac{\log n}{n}\Big)^{1/2}.$
@@
This is exactly the row-wise high-probability bound proved in the supplementary material of Zhang et al. (2017): after the decomposition (1), the row-averages of the two terms are bounded in (15) and (16) of their supplementary material, uniformly for all $i$ on an event of probability $1-O(n^{-c})$. ∎
Corollary 3.4 (Leave-one-out ZLZ row-wise risk and $(U_{ij},B_{ij})$ decomposition).
Fix $j\in[n]$. Let $A^{(-j)}$ be the adjacency matrix obtained by deleting row and column $j$ from $A$, and let $\widetilde{P}$ denote the ZLZ neighborhood-smoothing estimator computed from $A^{(-j)}$, with neighborhoods $N_{i}^{(-j)}$ defined exactly as in ZLZ but using the two-hop matrix $(A^{(-j)})^{2}/(n-1)$. Recall (3.1). For each $i\neq j$ and $\ell\neq j$, define
$U^{(-j)}_{i\ell}:=\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}(A_{k\ell}-P_{k\ell}),\qquad B^{(-j)}_{i\ell}:=\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}(P_{k\ell}-P_{i\ell}),$